Please show work, thank you in advance.

What is the coefficient of variation of the following income statement sales projection given the following information?

Possible Sales Level (in 000's)

Probability

1000

10%

2000

15%

2500

50%

2800

20%

3000

5%

  1. 1.41
  2. .22 (correct)
  3. 12%
  4. 18.2x

What is the standard deviation of the following income statement sales projection given the following information?

Possible Sales Level (in 000's)

Probability

0

10%

100

15%

400

40%

450

30%

600

5%

  1. $169.26 (correct)
  2. $340
  3. $387.50
  4. $400
Answer #1

Answer to Question 1:

Expected Value = 0.10 * 1,000 + 0.15 * 2,000 + 0.50 * 2,500 + 0.20 * 2,800 + 0.05 * 3,000
Expected Value = 2,360

Variance = 0.10 * (1,000 - 2,360)^2 + 0.15 * (2,000 - 2,360)^2 + 0.50 * (2,500 - 2,360)^2 + 0.20 * (2,800 - 2,360)^2 + 0.05 * (3,000 - 2,360)^2
Variance = 273,400

Standard Deviation = (273,400)^(1/2)
Standard Deviation = 522.87666

Coefficient of Variation = Standard Deviation / Expected Value
Coefficient of Variation = 522.87666 / 2,360
Coefficient of Variation = 0.22

Answer to Question 2:

Expected Value = 0.10 * 0 + 0.15 * 100 + 0.40 * 400 + 0.30 * 450 + 0.05 * 600
Expected Value = 340

Variance = 0.10 * (0 - 340)^2 + 0.15 * (100 - 340)^2 + 0.40 * (400 - 340)^2 + 0.30 * (450 - 340)^2 + 0.05 * (600 - 340)^2
Variance = 28,650

Standard Deviation = (28,650)^(1/2)
Standard Deviation = 169.26